# Explainable Metrics for the Assessment of Neurodegenerative Diseases through Handwriting Analysis

### v3

@@ Abstract
Motor dysfunction is a common sign of neurodegenerative diseases (NDs) such as Parkinson's disease (PD) and Alzheimer’s disease (AD), but may be difficult to detect, especially in the early stages. In this work, we examine the behavior of a wide array of explainable metrics extracted from the handwriting signals of 113 subjects performing multiple tasks on a digital tablet, as part of the Neurological Signals dataset. The aim is to measure their effectiveness in characterizing NDs, including AD and PD. To this end, task-agnostic and task-specific metrics are extracted from 14 distinct tasks. Subsequently, through statistical analysis and a series of classification experiments, we investigate which metrics provide greater discriminative power between NDs and healthy controls and amongst different NDs. Preliminary results indicate that the tasks at hand can all be effectively leveraged to distinguish between the considered set of NDs, specifically by measuring the stability, the speed of writing, the time spent not writing, and the pressure variations between groups from our handcrafted explainable metrics, which shows $p$-values lower than 0.0001 for multiple tasks. Using various binary classification algorithms on the computed metrics, we obtain up to 87% accuracy for the discrimination between AD and healthy controls (CTL), and up to 69% for the discrimination between PD and CTL.
Index Terms:
Handwriting Analysis, Alzheimer's Disease, Parkinson's Disease, Digital Biomarkers, Neurological Signals
@@ I Introduction
Numerous handwriting datasets have been introduced to study NDs, capturing either static images or dynamic time-series signals. However, most datasets are restricted to a single disease type and a single data modality, thereby limiting their experimental scope. For PD: HandPD [15]: Includes 92 participants (74 PD, 18 CTL), with image-based recordings of spiral and meander tracing. Naive Bayes classification achieved up to 78.9% accuracy. NewHandPD [15]: Expanded the previous dataset by adding pressure, acceleration, and tilt data from 14 PD and 21 CTL participants. Using CNNs, the study reached 83.77% accuracy on spirals and 87.14% on meanders. PaHaW [6]: Collected position and pressure data from spirals and repeated text in 37 PD and 38 CTL subjects. Using explainable metrics and SVMs, it achieved 81.3% accuracy. ParkinsonHW [16]: Recorded dynamic handwriting from 62 PD and 15 CTL subjects via a digital tablet. Cascaded non-neural classifiers achieved 84.67% accuracy on spirals and 90.91% on stability tasks [17].
@@
For AD: Early work by [14] evaluated handwriting signatures from 29 AD and 30 CTL subjects, achieving a 3% error rate with a decision tree model. The dataset introduced by [13] included 23 AD, 12 MCI, and 17 CTL participants performing drawing and writing tasks on a tablet. Sentence copying reached up to 80.4% classification accuracy across three groups. The DARWIN dataset [12], now a standard benchmark, includes 100 AD and 90 CTL participants engaged in copying, memory, and graphical tasks. Fusing non-neural classifiers trained on individual tasks yielded a top accuracy of 94.29%.
@@
Handwriting tasks for ND assessment can be categorized by complexity and task structure: Simple motor tasks involve drawing basic shapes such as spirals or straight lines, or stability exercises, such as maintaining a fixed pen over a tablet. These evaluate tremor, trajectory smoothness, velocity, and acceleration [18]. Written language tasks involve producing words or sentences. On tablets, these tasks enable the study of in-air movement and motor planning between words [18]. In-air time reflects cognitive preparation [5], and since in-air and on-surface dynamics are statistically independent [19], both provide complementary information. Increased in-air duration may reflect hesitation or planning deficits. Complex visuoconstructive tasks include copying figures or drawing clocks. The clock drawing test is an established tool for evaluating visuospatial and executive function across several NDs [20, 21]. Copying tasks allow examination of written response fidelity in varied contexts [22].
@@
In this work, we introduce a new dataset composed of digital handwriting samples from 103 participants, including individuals diagnosed with AD, Mild Cognitive Impairment (MCI), PD, and Related Disorders (a heterogeneous group we refer to as PDM, comprising atypical Parkinsonism and misdiagnosed cases). Participants completed 14 tablet-based handwriting tasks specifically designed to probe cognitive and motor deficits across these conditions. We extract a large set of explainable, quantitative metrics, referred to as handwriting metrics, from these tasks. These include both: Task-agnostic metrics, which capture global signal properties across all tasks, regardless of their nature, and Task-specific metrics, which target metrics tailored to particular task types. We then assess which metrics best distinguish between groups through statistical analyses and binary classification experiments. In particular, we compare ND subgroups (PD, AD, PDM) to healthy controls, and contrast PD and PDM to identify discriminative metrics within motor phenotypes.
@@
  • Design and evaluation of a broad set of task-agnostic and task-specific metrics that align with expected behavioral patterns and reveal significant group differences.
@@
The sections of the manuscripts are organized as follows: Section II presents the data collection pipeline and the main pre-processing stages. Section III introduces the task-agnostic metrics extracted from all tasks and the task-specific metrics designed for specific groups of tasks. Then, various methods used to gauge the significance of our metrics in the various tasks analyzed are described. Finally, we introduce the methods used for training classifiers for NDs assessment from the presented explainable metrics. Section IV shows and discusses our statistical and classification results. Section V contains the main findings, discussions, limits and conclusions of the study.
@@ II-A Data Collection
All participants signed informed consent, and the data collection was approved by the Johns Hopkins Medicine Institutional Review Board. Participants with PD continued their usual pharmacological treatment and took dopaminergic medication before the recording session. Handwriting samples were collected using a Wacom One 13 tablet [^1] https://estore.wacom.com/en-us/wacom-one-13-touch-dth134w0a.html with the associated pen. This setup enabled us to record the pressure applied by the pen on the tablet surface and its position both on the tablet and in the air when in close proximity. Note the recording session also included collection of saccadometric and synchronized speech, although these two modalities were not utilized in this study. Participants completed 14 written tasks per session, detailed in Section II-B, guided by a research assistant. We collected recordings for each participant for at least one session, while 21 participants were recorded up to three times, when available and interested, with each session spaced six months apart. The handwriting data collected is part of a larger, multimodal dataset that is yet to be published. The distribution of participants, with their age, gender, and total number of tasks recorded, is presented in Table I Participants classified as AD and MCI were combined into a single group, named AD in this study. Only a subset of the recorded participants were kept in that study, to balance as much as possible the age and gender of the participants. However, given the low number of female participants, gender balancing for the AD group was not possible, so we chose to keep the whole group.
@@ II-B Handwriting Tasks
Figure 1: Example of each type of task from all the four experimental groups. When on tablet, the color of the points transits from red to purple. When in air, the points are depicted in gray.
@@ II-B1 Point Tasks
The point tasks are designed to assess the static motor control and stability of the participants. Each participant was asked to maintain the pen vertically above a point shown in the tablet without moving or touching the tablet for $10\text{\,}\sec$. More specifically, in the Point Dominant and Point Non-Dominant tasks, participants are asked to use their dominant and Non-Dominant hand respectively, and in the Point Sustained, they are asked to use their dominant hand, in addition, to sustain the phonation of the vowel A during the exercise.
@@ III Methods
This section outlines the metrics calculated for each task and the methods used to assess their significance among different experimental groups. We developed interpretable metrics from raw handwriting signals collected during various tasks. Our analysis has two aspects: first, we examine dynamic handwriting properties common to all tasks; second, we evaluate performance in specific tasks, such as stability in point tasks and proximity between drawing and target in copying tasks.
III-A Metrics computation pipeline
To compute the metrics analyzed in this study, we follow the pipeline described in Figure 3. Each data file can be described as a time series containing $N\in\mathbb{R}$ points:
@@ III-A Metrics computation pipeline
We computed various metrics from the clean files, targeting prioritizing motor, cognitive or behavioral elements. The metrics can be categorized relative to or irrespective of the task:
@@
    Task-agnostic (see Section III-B): these metrics measure primarily dynamic properties of the signal such as speed, noise, and basic ratios.
@@
    Task-specific (see Section III-C): these metrics are designed for specific tasks or groups of tasks.
III-B Task-Agnostic Metrics
Task-agnostic metrics quantify the dynamic aspects of the drawings. They are computed using several signal-processing techniques. The extraction process, illustrated in Figure 3, may be parsed:
@@ III-B Task-Agnostic Metrics
    segmentation of each series using only the segments gathered In-Air or On-Tablet, and additionally decomposing the On-Tablet sequence in a suite of strokes. A stroke being defined as the trajectory On-Tablet between two In-Air moments. The segmentation is described in Section III-B2
@@
    computation of statistics (e.g., mean, the standard deviation) on the time-series obtained from the derivation of time independent metrics, described in Section III-B3.
Figure 3: Pipeline used to extract task-agnostic metrics across all tasks. The amplitudes for angular speech and acceleration are not kept, which is why we have 54 and not 58 metrics.
@@ III-B1 Time series
    Positions: horizontal and vertical speed and acceleration.
@@ III-B2 Segmentation
We implemented segmentation to distinguish between segments occurring on the tablet surface and those in the air. Subsequently, we computed time-independent metrics for the entire file, focusing either solely on points recorded on the tablet or those recorded in the air. Given that the point task primarily involves actions in the air, statistical computations are exclusively performed on the in-air segments. Conversely, for spiral tasks, only segments registered on the tablet are utilized. However, for writing and drawing tasks, both in-air and on-tablet segments are utilized. Notably, metrics based on pressure are not retained for the in-air segments, as they would be derived from a constant value.
III-B3 Time-Independent Metrics
@@
We chose not to utilize the amplitude measure for angular velocity and accelerations, as it would always yield identical values. Furthermore, we calculated the Shannon Entropy based on the points' positions, whether on the tablet or in the air. To achieve this, we employ a Gaussian kernel density estimate (KDE) on the points' $X$ and $Y$ positions. This enabled us to determine the probability $q_{i}$ for each point $(x_{i},y_{i})$. Shannon Entropy is then defined as $H_{s}=-\sum_{i=1}^{N}q_{i}ln(q_{i})$. Following the computation of these metrics, we obtained a total of 54 task-agnostic metrics across the 14 tasks.
III-C Task-Specific Metrics
In this section, we detail the various task-specific metrics we computed. We divided the presentations of the metrics along the four main tasks' categories proposed in Section II-B, with the addition of a section specific to the Numbers task.
III-C1 Point Tasks Metrics
This section presents the metrics specific to the Point Tasks. Figure 4 depicts these metrics. Point Tasks are mainly used to measure the motor stability of participants and are mostly used for characterization of tremor, primarily relevant to PD or PDM (though tremor may also be observed in AD, due to copathology). Most dynamic metrics have already been explored as task-agnostic metrics, so we focused on the static metrics of the drawing: how symmetric and regular it is, and what the dispersion of the samples is. For this task, we calculated the geometric center of the pen's position over the digital tablet. We used this geometric center as a pivot to compare the distribution of the points with respect to the computed center, through 3 metrics:
@@ III-C1 Point Tasks Metrics
Figure 4: Schematic of the characteristics measured by different point-tasks specific metrics.
III-C2 Spiral Tasks Metrics
This section discusses static and dynamic metrics used to analyze hand-drawn spirals.
Those metrics are following three directions, illustrated in Figure 5:
@@ III-C2 Spiral Tasks Metrics
Figure 5: Spiral-specific metrics computation process
@@
As shown in Figure 5, the participants produce rounded spirals for all the spiral tasks. Our approach involves initially characterizing the properties of the outer boundaries of the spiral as an ellipse, leveraging these properties as metrics. We used Principal Component Analysis (PCA) to find the directions of the semi-major and semi-minor axes of each spiral, then find the points further away from the geometric center of the spiral in both of those directions to obtain a measure of the semi-minor axis $a$ and the semi-major axis $b$. Using those values, we define a set of metrics:
@@
Once the outside shape has been characterized, we looked into the properties of the arms of the spiral. Each spiral is split into segments, containing each one full loop. For each loop, we computed its radius as the average distances between its points and the geometric center of the spiral, which allowed us to compute another set of metrics:
@@
  • The standard deviation of the loops radius differences: the standard deviation of the radius difference between consecutive loops. This metric quantifies the regularity of the spiral.
@@
Dynamic metrics are then computed based on the total time taken to complete the task:
@@
III-C3 Numbers Task Metrics
This section focuses on the metrics unique to the Numbers task. Participants engaged in solving addition problems by moving down the tablet, allowing for individual analysis of each response. This methodology enabled us to track the evolution of handwriting metrics as participants navigate through the task. For a comprehensive breakdown of these metrics, please refer to Figure 6 below.
Figure 6: Schematic of the metrics used for each answer of the Numbers task. The left graph shows the variations in pressures across the whole task, and the right schematic shows the metrics measured between 2 consecutive answers.

### v4

# Interpretable Features for the Assessment of Neurodegenerative Diseases through Handwriting Analysis

@@ Abstract
Motor dysfunction is a common sign of neurodegenerative diseases (NDs) such as Parkinson's disease (PD) and Alzheimer’s disease (AD), but may be difficult to detect, especially in the early stages. In this work, we examine the behavior of a wide array of interpretable features extracted from the handwriting signals of 113 subjects performing multiple tasks on a digital tablet, as part of the Neurological Signals dataset. The aim is to measure their effectiveness in characterizing NDs, including AD and PD. To this end, task-agnostic and task-specific features are extracted from 14 distinct tasks. Subsequently, through statistical analysis and a series of classification experiments, we investigate which features provide greater discriminative power between NDs and healthy controls and amongst different NDs. Preliminary results indicate that the tasks at hand can all be effectively leveraged to distinguish between the considered set of NDs, specifically by measuring the stability, the speed of writing, the time spent not writing, and the pressure variations between groups from our handcrafted interpretable features, which shows a statistically significant difference between groups, across multiple tasks. Using various binary classification algorithms on the computed features, we obtain up to 87% accuracy for the discrimination between AD and healthy controls (CTL), and up to 69% for the discrimination between PD and CTL.
@@ I Introduction
Numerous handwriting datasets have been introduced to study NDs, capturing either static images or dynamic time-series signals. However, most datasets are restricted to a single disease type and a single data modality, thereby limiting their experimental scope. For the assessment of PD:
  • HandPD [15]: Includes 92 participants (74 PD, 18 CTL), with image-based recordings of spiral and meander tracing. Naive Bayes classification achieved up to 78.9% accuracy.
  • NewHandPD [16]: Expanded the previous dataset by adding pressure, acceleration, and tilt data from 14 PD and 21 CTL participants. Using CNNs, the study reached 83.77% accuracy on spirals and 87.14% on meanders.
  • PaHaW [6]: Collected position and pressure data from spirals and repeated text in 37 PD and 38 CTL subjects. Using interpretable features and SVMs, it achieved 81.3% accuracy.
  • ParkinsonHW [17]: Recorded dynamic handwriting from 27 PD and 28 CTL subjects via a digital tablet. Cascaded non-neural classifiers achieved 84.67% accuracy on spirals and 90.91% on stability tasks [18].
For the assessment of AD:
  • Early work by [14] evaluated handwriting signatures from 29 AD and 30 CTL subjects, achieving a 3% error rate with a decision tree model.
  • The dataset introduced by [13] included 23 AD, 12 MCI, and 17 CTL participants performing drawing and writing tasks on a tablet. Sentence copying reached up to 80.4% classification accuracy across three groups.
@@
  • The DARWIN dataset [12], now a standard benchmark, includes 89 AD and 85 CTL participants engaged in copying, memory, and graphical tasks. Fusing non-neural classifiers trained on individual tasks yielded a top accuracy of 94.29%.
@@
Handwriting tasks for ND assessment can be categorized by complexity and task structure: Simple motor tasks involve drawing basic shapes such as spirals or straight lines, or stability exercises, such as maintaining a fixed pen over a tablet. These evaluate tremor, trajectory smoothness, speed, and acceleration [19]. Written language tasks involve producing words or sentences. On tablets, these tasks enable the study of in-air movement and motor planning between words [19]. In-air time reflects cognitive preparation [5], and since in-air and on-surface dynamics are non-redundant [20], both provide complementary information. Increased in-air duration may reflect hesitation or planning deficits. Complex visuoconstructive tasks include copying figures or drawing clocks. The clock drawing test is an established tool for evaluating visuospatial and executive function across several NDs [21, 22]. Copying tasks allow examination of written response fidelity in varied contexts [23].
@@
In this work, we introduce a new dataset composed of digital handwriting samples from 103 participants, including individuals diagnosed with AD, Mild Cognitive Impairment (MCI), PD, and Related Disorders (a heterogeneous group we refer to as PDM, comprising atypical Parkinsonism and misdiagnosed cases). Participants completed 14 tablet-based handwriting tasks specifically designed to probe cognitive and motor deficits across these conditions. We extract a large set of interpretable, quantitative measurements, referred to as handwriting features, from these tasks. These include both: Task-agnostic features, which capture global signal properties across all tasks, regardless of their nature, and Task-specific features, which target features tailored to particular task types. We then assess which features best distinguish between groups through statistical analyses and binary classification experiments. In particular, we compare ND subgroups (PD, AD, PDM) to healthy controls, and contrast PD and PDM to identify discriminative features within motor phenotypes.
@@
  • Design and evaluation of a broad set of task-agnostic and task-specific features that align with expected behavioral patterns and reveal significant group differences.
@@
The sections of the manuscripts are organized as follows: Section II presents the data collection pipeline and the main pre-processing stages. Section III introduces the task-agnostic features extracted from all tasks and the task-specific features designed for specific groups of tasks. Then, various methods used to gauge the significance of our features in the various tasks analyzed are described. Finally, we introduce the methods used for training classifiers for NDs assessment from the presented interpretable features. Section IV shows and discusses our statistical and classification results. Section V contains the main findings, discussions, limits and conclusions of the study.
@@ II-A Data Collection
All participants signed informed consent, and the data collection was approved by the Johns Hopkins Medicine Institutional Review Board. Participants with PD continued their usual pharmacological treatment and took dopaminergic medication before the recording session. Handwriting samples were collected using a Wacom One 13 tablet [^1] https://estore.wacom.com/en-us/wacom-one-13-touch-dth134w0a.html with the associated pen. This setup enabled us to record the pressure applied by the pen on the tablet surface and its position both on the tablet and in the air when in close proximity. Note the recording session also included collection of saccadometric and synchronized speech, although these two modalities were not utilized in this study. Participants completed 14 written tasks per session, detailed in Section II-B, guided by a research assistant. We collected recordings for each participant for at least one session, while 21 participants were recorded up to three times, when available and interested, with each session spaced six months apart. The handwriting data collected is part of a larger, multimodal dataset that is yet to be published. The distribution of participants, with their age, gender, and total number of tasks recorded, is presented in Table I. Participants classified as AD and MCI were combined into a single group, named AD* in this study. Only a subset of the recorded participants were kept in that study, to balance as much as possible the age and gender of the participants. However, given the low number of female participants, gender balancing for the AD* group was not possible, so we chose to keep the whole group.
@@ II-B Handwriting Tasks
Figure 1: Example of each type of task from all four experimental groups. When on tablet, the color of the points transits from red to purple. When in air, the points are depicted in gray.
@@ II-B1 Point Tasks
The point tasks are designed to assess the static motor control and stability of the participants. Each participant was asked to maintain the pen vertically above a point shown in the tablet without moving or touching the tablet for $10\text{\,}\sec$. More specifically, in the Point Dominant and Point Non-Dominant tasks, participants are asked to use their dominant and Non-Dominant hand, respectively, and in the Point Sustained, they are asked to use their dominant hand, in addition, to sustain the phonation of the vowel A during the exercise.
@@ III Methods
This section outlines the features calculated for each task and the methods used to assess their significance among different experimental groups. We developed interpretable features from raw handwriting signals collected during various tasks. Our analysis has two aspects: first, we examine dynamic handwriting properties common to all tasks; second, we evaluate performance in specific tasks, such as stability in point tasks and proximity between drawing and target in copying tasks.
III-A Features computation pipeline
To compute the features analyzed in this study, we follow the pipeline described in Figure 3. Each data file can be described as a time series containing $N\in\mathbb{R}$ points:
@@ III-A Features computation pipeline
We computed various features from the clean files, targeting prioritizing motor, cognitive or behavioral elements. The features can be categorized relative to or irrespective of the task:
@@
    Task-agnostic (see Section III-B): these features measure primarily dynamic properties of the signal such as speed, noise, and basic ratios.
@@
    Task-specific (see Section III-C): these features are designed for specific tasks or groups of tasks.
III-B Task-Agnostic Features
Task-agnostic features quantify the dynamic aspects of the drawings. They are computed using several signal-processing techniques. The extraction process, illustrated in Figure 3, may be parsed:
@@ III-B Task-Agnostic Features
    segmentation of each series using only the segments gathered In-Air or On-Tablet. A stroke being defined as the trajectory On-Tablet between two In-Air moments. The segmentation is described in Section III-B2
@@
    computation of statistics (e.g., mean, the standard deviation) on the time-series obtained from the derivation of time independent features, described in Section III-B3.
Figure 3: Pipeline used to extract task-agnostic features across all tasks. From the raw data, 5 time series are extracted, differentiated twice, and then the segmentation selects only the in-air parts (yellow path) or the on-tablet parts (brown path), and statistics and entropy are computed from each segmented time-series. The amplitudes for angular velocity and acceleration are not kept, which is why we have 54 and not 58 features.
@@ III-B1 Time series
    Positions: horizontal and vertical velocity and acceleration.
@@ III-B2 Segmentation
We implemented segmentation to distinguish between segments occurring on the tablet surface and those in the air. Subsequently, we computed time-independent features for the entire file, focusing either solely on points recorded on the tablet or those recorded in the air. Given that the point task primarily involves actions in the air, statistical computations are exclusively performed on the in-air segments. Conversely, for spiral tasks, only segments registered on the tablet are utilized. However, for writing and drawing tasks, both in-air and on-tablet segments are utilized. Notably, features based on pressure are not retained for the in-air segments, as they would be derived from a constant value.
III-B3 Time-Independent Features
@@
We chose not to utilize the amplitude measure for angular velocity and acceleration, as it would always yield identical values. Furthermore, we calculated the Shannon Entropy based on the points' positions, whether on the tablet or in the air. To achieve this, we employ a Gaussian kernel density estimate (KDE) on the points' $X$ and $Y$ positions. This enabled us to determine the probability $q_{i}$ for each point $(x_{i},y_{i})$. Shannon Entropy is then defined as $H_{s}=-\sum_{i=1}^{N}q_{i}ln(q_{i})$. Following the computation of these features, we obtained a total of 54 task-agnostic features across the 14 tasks.
III-C Task-Specific Features
In this section, we detail the various task-specific features we computed. We divided the presentations of the features along the four main tasks' categories proposed in Section II-B, with the addition of a section specific to the Numbers task.
III-C1 Point Tasks Features
This section presents the features specific to the Point Tasks. Figure 4 depicts these features. Point Tasks are mainly used to measure the motor stability of participants and are mostly used for characterization of tremor, primarily relevant to PD or PDM (though tremor may also be observed in AD, due to copathology). Most dynamic features have already been explored as task-agnostic features, so we focused on the static features of the drawing: how symmetric and regular it is, and what the dispersion of the samples is. For this task, we calculated the geometric center of the pen's position over the digital tablet. We used this geometric center as a pivot to compare the distribution of the points with respect to the computed center, through 3 features:
@@ III-C1 Point Tasks Features
Figure 4: Schematic of the characteristics measured by different point-tasks specific features.
III-C2 Spiral Tasks Features
This section discusses static and dynamic features used to analyze hand-drawn spirals.
Those features are following three directions, illustrated in Figure 5:
@@ III-C2 Spiral Tasks Features
Figure 5: Spiral-specific features computation process
@@
As shown in Figure 5, the participants produce rounded spirals for all the spiral tasks. Our approach involves initially characterizing the properties of the outer boundaries of the spiral as an ellipse, leveraging these properties as features. We used Principal Component Analysis (PCA) to find the directions of the semi-major and semi-minor axes of each spiral, then find the points further away from the geometric center of the spiral in both of those directions to obtain a measure of the semi-minor axis $a$ and the semi-major axis $b$. Using those values, we define a set of features:
@@
Once the outside shape has been characterized, we looked into the properties of the arms of the spiral. Each spiral is split into segments, containing each one full loop. For each loop, we computed its radius as the average distances between its points and the geometric center of the spiral, which allowed us to compute another set of features:
@@
  • The standard deviation of the loops radius differences: the standard deviation of the radius difference between consecutive loops. This feature quantifies the regularity of the spiral.
@@
Dynamic features are then computed based on the total time taken to complete the task:
@@
III-C3 Numbers Task Features
This section focuses on the features unique to the Numbers task. Participants engaged in solving addition problems by moving down the tablet, allowing for individual analysis of each response. This methodology enabled us to track the evolution of handwriting features as participants navigate through the task. For a comprehensive breakdown of these features, please refer to Figure 6 below.
Figure 6: Schematic of the features used for each answer of the Numbers task. The left graph shows the variations in pressures across the whole task, and the right schematic shows the features measured between 2 consecutive answers.